# Joint Design of Radar Receive Filter and Unimodular ISAC Waveform with Sidelobe Level Control

### v1

@@ III-C3 Solving for Transmit Waveform
[…]
whose closed form solution is
@@ III-D Convergence Analysis
Before presenting the convergence results of the proposed Algorithm to the stationary point of problem (8a), we first formally define its stationary point. The idea is to first transform problem (8a) into an equivalent problem defined in the real space and then use the classic definition in the real space. Specifically, by denoting $\tilde{\mathbf{f}}=[\mathcal{R}\{\text{vec}\{\mathbf{F}^{\text{H}}\}\}^{\text{T}},\mathcal{I}\{\text{vec}\{\mathbf{F}^{\text{H}}\}\}^{\text{T}}]^{\text{T}}$, $\mathcal{S}_{\tilde{\mathbf{f}}}=\{\tilde{\mathbf{f}}\in\mathbb{R}^{2N_{T}T\times 1}\mid\|\tilde{\mathbf{f}}\|_{2}^{2}=P_{\text{F}}\}$, $\mathcal{S}_{\tilde{\mathbf{c}}}=\{\tilde{\mathbf{c}}\in\mathbb{R}^{2N_{\text{T}}^{2}\times 1}\mid\|\tilde{\mathbf{c}}\|_{2}\leq\xi\}$, and $\mathbf{D}_{k}=\mathbf{I}\otimes(\mathbf{X}\mathbf{J}_{k})$ with $\tilde{\mathbf{D}}_{k}=\left[\begin{smallmatrix}\mathcal{R}\{\mathbf{D}_{k}\}&-\mathcal{I}\{\mathbf{D}_{k}\}\\ \mathcal{I}\{\mathbf{D}_{k}\}&\mathcal{R}\{\mathbf{D}_{k}\}\end{smallmatrix}\right]$, problem (8a) w.r.t. $\mathbf{F}$ is equivalent to the following real-valued vector minimization problem:
$\min_{\tilde{\mathbf{f}}}g(\tilde{\mathbf{f}})+\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{f}}}}(\tilde{\mathbf{f}})+\sum_{k\in\mathbf{\Omega}_{K}}\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{c}}}}(\tilde{\mathbf{D}}_{k}\tilde{\mathbf{f}}),$ (45)
where $g(\tilde{\mathbf{f}})$ denotes the reformulated objective function $f(\mathbf{F},\mathbf{X})$ in problem (8a) w.r.t. the real-valued vector $\tilde{\mathbf{f}}$. Similarly, by denoting $\tilde{\mathbf{x}}=[\mathcal{R}\{\text{vec}\{\mathbf{X}\}\}^{\text{T}},\mathcal{I}\{\text{vec}\{\mathbf{X}\}\}^{\text{T}}]^{\text{T}}$,
$\mathcal{S}_{\tilde{\mathbf{x}}}$ =$=\{\tilde{\mathbf{x}}\in\mathbb{R}^{2N_{T}T\times 1}\mid\|\tilde{\mathbf{x}}_{[i]}\|_{2}=P_{i},\|\tilde{\mathbf{x}}_{[i]}\|_{\infty}\leq\eta^{\prime},~{}\forall~{}i\in[1,N_{\text{T}}]\},$
$\tilde{\mathbf{x}}_{[i]}$ $=\left[[\tilde{x}_{i},\dots,\tilde{x}_{TN_{\text{T}}+i}]^{\text{T}},[\tilde{x}_{(T+1)N_{\text{T}}+i},\dots,\tilde{x}_{(2T-1)N_{\text{T}}+i}]^{\text{T}}\right]^{\text{T}},$
and $\mathbf{E}_{k}=(\mathbf{J}_{k}\mathbf{F}^{\text{H}})^{\text{T}}\otimes\mathbf{I}$ with $\tilde{\mathbf{E}}_{k}=\left[\begin{smallmatrix}\mathcal{R}\{\mathbf{E}_{k}\}&-\mathcal{I}\{\mathbf{E}_{k}\}\\ \mathcal{I}\{\mathbf{E}_{k}\}&\mathcal{R}\{\mathbf{E}_{k}\}\end{smallmatrix}\right]$, problem (8a) w.r.t. $\mathbf{X}$ is equivalent to
$\min_{\tilde{\mathbf{x}}}h(\tilde{\mathbf{x}})+\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{x}}}}(\tilde{\mathbf{x}})+\sum_{k\in\mathbf{\Omega}_{K}}\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{c}}}}(\tilde{\mathbf{E}}_{k}\tilde{\mathbf{x}}),$ (46)
where $h(\tilde{\mathbf{x}})$ denotes the reformulated objective function in problem (8a) w.r.t. the real-valued vector $\tilde{\mathbf{x}}$.
Based on the equivalence between problem (8a) w.r.t. $\mathbf{F}$ and $\mathbf{X}$ and problems (45) and (46), we have the following definition of the stationary point of problem (8a).
Definition 2 (Stationary point of problem (8a)).
The point $(\bar{\mathbf{F}},\bar{\mathbf{X}})$ is a stationary point of problem (8a), if for $\bar{\mathbf{f}}=[\mathcal{R}\{{\rm vec}\{\bar{\mathbf{F}}^{{\rm H}}\}\}^{{\rm T}},\allowbreak\mathcal{I}\{{\rm vec}\{\bar{\mathbf{F}}^{{\rm H}}\}\}^{{\rm T}}]^{{\rm T}}$, $\bar{\mathbf{x}}=[\mathcal{R}\{{\rm vec}\{\bar{\mathbf{X}}\}\}^{{\rm T}},\mathcal{I}\{{\rm vec}\{\bar{\mathbf{X}}\}\}^{{\rm T}}]^{{\rm T}}$, there exist $\bar{\mathbf{a}}\in\partial\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{f}}}}(\bar{\mathbf{f}})$, $\bar{\mathbf{b}}\in\partial\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{x}}}}(\bar{\mathbf{x}})$, and $\bar{\mathbf{p}}_{k}\in\partial\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{c}}}}(\tilde{\mathbf{D}}_{k}\bar{\mathbf{f}})$ and $\mathbf{q}_{k}\in\partial\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{c}}}}(\tilde{\mathbf{E}}_{k}\bar{\mathbf{x}})$ for all $k\in\mathbf{\Omega}_{K}$, such that
$\mathbf{0}$ $=\bar{\mathbf{a}}+\nabla_{\tilde{\mathbf{f}}}g(\bar{\mathbf{f}})+\sum_{k\in\mathbf{\Omega}_{K}}\tilde{\mathbf{D}}_{k}^{\text{T}}\bar{\mathbf{p}}_{k},$
$\mathbf{0}$ $=\bar{\mathbf{b}}+\nabla_{\tilde{\mathbf{x}}}h(\bar{\mathbf{x}})+\sum_{k\in\mathbf{\Omega}_{K}}\tilde{\mathbf{E}}_{k}^{\text{T}}\bar{\mathbf{q}}_{k}.$
Now we can present our main theoretical result.
@@
Any limit point of $\{(\mathbf{F}^{(\ell+1)},\mathbf{X}^{(\ell+1)})\}_{\ell\geq 0}$ generated by Algorithm 1 is a stationary point of problem (8a).
@@ IV Numerical Results
In this section, we present simulation results to evaluate the performance of the proposed algorithms under various parameter configurations. The communication channel $\mathbf{H}$ is generated by following [27] with Extended Pedestrian A (EPA) fading profile [9]. The information symbol matrix $\mathbf{S}$ is modulated using a unit-power QPSK alphabet, with each entry in $\mathbf{S}$ having unit power. The desired radar receive spatial response, $\mathbf{R}_{\text{d}}$, is considered to be a $3$ dB beamwidth of $20^{\circ}$ focusing at $0^{\circ}$, generated by the algorithm in [34].
@@
$K$ Maximum index of range bin of interest 6
@@
$P_{\text{F}}$ Power of the receive filter 64
@@ IV-A Convergence Performance
(b) The sum rate under SNR = $6$ dB and beampattern MSE w.r.t $K$.
@@ IV-B Communication and Radar Performance under Different System Configurations
In this subsection, we evaluate the communication and radar performance of the proposed algorithm under different system configurations. Specifically, we examine the impact of the range sidelobe suppression area size determined by $K$, and the maximum sidelobe level constraints $\xi^{\prime}$ (dB).
Fig. 5 illustrates the impact of the sidelobe suppression area sizes determined by $K$ on the optimized results. As shown in Fig. 5LABEL:sub@fig:SL_K, even with the increment of $K$, the optimized sidelobe levels can still approach the predetermined threshold. In Fig. 5LABEL:sub@fig:rate_MSE_K, the achievable sum rate decreases and the beampattern MSE increases as $K$ grows, since a larger suppression area imposes more constraints, reducing the feasible region and degrading the performance of optimized results. This higher beampattern MSE leads to a slight decrease in the mainlobe level.
@@
(b) The sum rate under SNR = $6$ dB and beampattern MSE w.r.t. PSLR constraints.
@@
Fig. 6 shows the impact of maximum sidelobe level constraints on the optimized results. As seen in Fig. 6LABEL:sub@fig:SL_PSLR, the sidelobe constraints can always be met. Lower sidelobe constraints require $\mathbf{X}$ and $\mathbf{F}$ to be less correlated, making the problem more complex and resulting in worse performance in both S&\&&C. This can be observed in Fig. 6LABEL:sub@fig:rate_MSE_PSLR, in which a lower maximum sidelobe level constraint leads to a worse achievable sum rate and a larger beampattern MSE.
@@ IV-C Comparison with Existing SOTA Approaches
Figure 8: The MIMO radar image under the results of different algorithms when SNR = $15$ dB: (a) Original image of “T”. (b) The image formed by the proposed ALM algorithm. (c) The image formed by the Exact Design scheme. (d) The image formed by the fixed penalty ALM algorithm.
@@
In this subsection, we compare the radar and communication performance of the proposed approach with the modified work [11] and the ALM algorithm with fixed penalty parameter. In [11], the joint design of the radar receive filter and transmit waveform focuses on PSL control for pure radar sensing. We adapt the algorithm in [11] by first solving $\min_{\mathbf{X}\in\mathcal{S}_{\text{X}}}\|\mathbf{Hx}-\mathbf{s}\|_{2}^{2}$ using the BSUM method and then following the approach in [11] to obtain the radar receive filter $\mathbf{F}$. We refer to the modified algorithm of [11] as the “Exact Design” scheme. Furthermore, we also implemented the BSUM-based ALM algorithm with fixed penalty parameter to show the superiority of the proposed adaptive penalty parameter scheme. We refer to this as “Fixed Penalty” scheme. Without loss of generality, the penalty parameters are fixed to be the final value of the proposed ALM algorithm with adaptive penalty parameter scheme. In the legends of the result figures, “Direc” refers to the receive beampattern focused at $0^{\circ}$ with a $3$ dB beamwidth of $20^{\circ}$, generated by the algorithm in [34], and “Omni” denotes the omnidirectional desired beampattern, i.e., $\mathbf{R}_{\text{d}}=(P_{\text{X}}/N_{T})\mathbf{I}_{N_{T}}$.
@@
Fig. 7 presents the optimized results of different algorithms. In Fig. 7LABEL:sub@fig:rate_algs, the proposed ALM algorithm achieves the highest sum rate under the directional beampattern case. For an omnidirectional beam, we need to set a larger $\alpha$ to make the final beam pattern closer to omnidirectional because an omnidirectional beam requires that $\mathbf{X}$ and $\mathbf{F}$ be orthogonal in each row and column, which significantly shrinks the feasible region. However, this degrades the minimization on $P_{\text{MUI}}$ and lead to a lower rate. Fig. 7LABEL:sub@fig:bpMSE_algs shows that the Exact Design scheme perfectly matches the desired beampatterns in both the directional and omnidirectional cases while the two ALM algorithms have slight differences from the omnidirectional beampattern. Fig. 7LABEL:sub@fig:SL_algs shows the proposed ALM algorithm and ALM algorithm with fixed penalty parameter have similar sidelobe control performance, but the latter one has a much lower rate due to the inappropriate penalty parameter.
Fig. 8 shows the MIMO radar images formed by the results of different algorithms. Both the proposed ALM algorithm and the ALM algorithm with fixed penalty parameter achieve similar imaging performance but the achievable sum rate is much lower under the fixed penalty parameter scheme. Although the Exact Design scheme meets the desired beampattern perfectly, it has a high sidelobe level, which results in a poor MIMO radar image.
@@ Appendix D Proof of Theorem 2
The convergence proof consists of two main steps, where the first step shows the feasibility of the limit point of the sequence generated by Algorithm 1 and the second step shows that any limit point is a feasible stationary point of problem (8a). We present these two steps in two subsections separately.
@@ D-A Feasibility of the Limit Point
D-B Stationarity of the Limit Point
By reformulating problem (13) w.r.t. $\mathbf{F}$, $\mathbf{X}$, and $\bm{C}$ into real-valued vectors $\tilde{\mathbf{f}}$, $\tilde{\mathbf{x}}$, and $\tilde{\mathbf{c}}_{k}=[\mathcal{R}\{\mathbf{c}_{k}\}^{\text{T}},\mathcal{I}\{\mathbf{c}_{k}\}^{\text{T}}]^{\text{T}}$ with $\mathbf{c}_{k}=\text{vec}\{\mathbf{C}_{k}\}$ for all $k\in\mathbf{\Omega}_{K}$, and $\tilde{\mathbf{u}}_{k}=[\mathcal{R}\{\mathbf{u}_{k}\}^{\text{T}},\mathcal{I}\{\mathbf{u}_{k}\}^{\text{T}}]^{\text{T}}$ with $\mathbf{u}_{k}=\text{vec}\{\mathbf{U}_{k}\}$ for all $k\in\mathbf{\Omega}_{K}$, we have the following equivalent real-valued vector problems:
$\min_{\tilde{\mathbf{f}}}g(\tilde{\mathbf{f}})+\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{f}}}}(\tilde{\mathbf{f}})+\sum_{k\in\mathbf{\Omega}_{K}}\left[\tilde{\mathbf{u}}_{k}^{\text{T}}(\tilde{\mathbf{D}}_{k}\tilde{\mathbf{f}}-\tilde{\mathbf{c}}_{k})+\frac{\rho}{2}\|\tilde{\mathbf{D}}_{k}\tilde{\mathbf{f}}-\tilde{\mathbf{c}}_{k}\|_{2}^{2}\right],$
$\min_{\tilde{\mathbf{x}}}h(\tilde{\mathbf{x}})+\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{x}}}}(\tilde{\mathbf{x}})+\sum_{k\in\mathbf{\Omega}_{K}}\left[\tilde{\mathbf{u}}_{k}^{\text{T}}(\tilde{\mathbf{E}}_{k}\tilde{\mathbf{x}}-\tilde{\mathbf{c}}_{k})+\frac{\rho}{2}\|\tilde{\mathbf{E}}_{k}\tilde{\mathbf{x}}-\tilde{\mathbf{c}}_{k}\|_{2}^{2}\right],$
$\min_{\tilde{\mathbf{c}}_{k}}\mathbb{I}_{\tilde{\mathbf{c}}}(\tilde{\mathbf{c}}_{k})+\tilde{\mathbf{u}}_{k}^{\text{T}}(\tilde{\mathbf{v}}_{k}-\tilde{\mathbf{c}}_{k})+\frac{\rho}{2}\|\tilde{\mathbf{v}}_{k}-\tilde{\mathbf{c}}_{k}\|_{2}^{2},$
where $g(\tilde{\mathbf{f}})$ and $h(\tilde{\mathbf{x}})$ are defined in (45) and (46), respectively, $\tilde{\mathbf{v}}_{k}=[\mathcal{R}\{\mathbf{v}_{k}\}^{\text{T}},\mathcal{I}\{\mathbf{v}_{k}\}^{\text{T}}]^{\text{T}}$ with $\mathbf{v}_{k}=\text{vec}\{\mathbf{X}\mathbf{J}_{k}\mathbf{F}^{\text{H}}\}$. It is simple to verify that $\tilde{\mathbf{v}}_{k}=\tilde{\mathbf{D}}_{k}\tilde{\mathbf{f}}=\tilde{\mathbf{E}}_{k}\tilde{\mathbf{x}}$. Then, from the definition of the $\varepsilon$-stationary point of (13), we immediately get the following result: If the point $(\bar{\bm{C}},\bar{\mathbf{F}},\bar{\mathbf{X}})$ is an $\varepsilon$-stationary point of problem (13), then for $\bar{\mathbf{f}}=[\mathcal{R}\{{\rm vec}\{\bar{\mathbf{F}}^{{\rm H}}\}\}^{{\rm T}},\mathcal{I}\{{\rm vec}\{\bar{\mathbf{F}}^{{\rm H}}\}\}^{{\rm T}}]^{{\rm T}}$, $\bar{\mathbf{x}}=[\mathcal{R}\{{\rm vec}\{\bar{\mathbf{X}}\}\}^{{\rm T}},\mathcal{I}\{{\rm vec}\{\bar{\mathbf{X}}\}\}^{{\rm T}}]^{{\rm T}}$ and $\bar{\mathbf{c}}_{k}=[\mathcal{R}\{{\rm vec}\{\bar{\mathbf{C}}_{k}\}\}^{{\rm T}},\allowbreak\mathcal{I}\{{\rm vec}\{\bar{\mathbf{C}}_{k}\}\}^{{\rm T}}]^{{\rm T}}$ for all $k\in\mathbf{\Omega}_{K}$, there exist $\bar{\mathbf{a}}\in\partial\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{f}}}}(\mathbf{f})$, $\bar{\mathbf{b}}\in\partial\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{x}}}}(\bar{\mathbf{x}})$, $\bar{\mathbf{w}}_{k}\in\partial\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{c}}}}(\bar{\mathbf{c}}_{k})$, and $\{\bar{\mathbf{u}}_{k}\}_{k\in\mathbf{\Omega}_{K}}$, such that
$\|\bar{\mathbf{a}}+\nabla_{\tilde{\mathbf{f}}}g(\bar{\mathbf{f}},\bar{\mathbf{x}})+\sum_{k\in\mathbf{\Omega}_{K}}\tilde{\mathbf{D}}_{k}^{{\rm T}}[\bar{\mathbf{u}}_{k}+\rho(\tilde{\mathbf{D}}_{k}\bar{\mathbf{f}}-\bar{\mathbf{c}}_{k})]\|_{2}\leq\varepsilon,\\ \|\bar{\mathbf{b}}+\nabla_{\tilde{\mathbf{x}}}h(\bar{\mathbf{f}},\bar{\mathbf{x}})+\sum_{k\in\mathbf{\Omega}_{K}}\tilde{\mathbf{E}}_{k}^{{\rm T}}[\bar{\mathbf{u}}_{k}+\rho(\tilde{\mathbf{E}}_{k}\bar{\mathbf{x}}-\bar{\mathbf{c}}_{k})]\|_{2}\leq\varepsilon,\\ \|\bar{\mathbf{w}}_{k}-[\bar{\mathbf{u}}_{k}+\rho(\bar{\mathbf{v}}_{k}-\bar{\mathbf{c}}_{k})]\|_{2}\leq\varepsilon,~{}\forall~{}k\in\mathbf{\Omega}_{K}.$ (67)
From the results in Lemma 3 and Theorem 1, Algorithm 2 returns an $\varepsilon^{(\ell)}$-stationary point after a finite number of (inner) iterations. Taking $\ell\rightarrow\infty$ with $\varepsilon^{(\ell)}\rightarrow 0$ yields
$\mathbf{0}$ $=\mathbf{a}^{(\infty)}+\nabla_{\tilde{\mathbf{f}}}g(\tilde{\mathbf{f}}^{(\infty)})+\sum_{k\in\mathbf{\Omega}_{K}}\tilde{\mathbf{D}}_{k}^{{\rm T}}\mathbf{w}_{k}^{(\infty)},$
$\mathbf{0}$ $=\mathbf{b}^{(\infty)}+\nabla_{\tilde{\mathbf{x}}}h(\tilde{\mathbf{x}}^{(\infty)})+\sum_{k\in\mathbf{\Omega}_{K}}\tilde{\mathbf{E}}_{k}^{{\rm T}}\mathbf{w}_{k}^{(\infty)}.$
Together with the feasibility proved in the previous subsection, we can verify that $\tilde{\mathbf{v}}_{k}^{(\infty)}=\tilde{\mathbf{c}}_{k}^{(\infty)}$ for all $k\in\mathbf{\Omega}_{K}$. According to the limiting subdifferential defined in [29], we have $\mathbf{w}_{k}^{(\infty)}\in\partial\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{c}}}}(\tilde{\mathbf{D}}_{k}\tilde{\mathbf{f}}^{(\infty)})$ and $\mathbf{w}_{k}^{(\infty)}\in\partial\mathbb{I}_{\mathcal{S}_{\tilde{\mathbf{c}}}}(\tilde{\mathbf{E}}_{k}\tilde{\mathbf{x}}^{(\infty)})$ since $\tilde{\mathbf{v}}_{k}=\tilde{\mathbf{D}}_{k}\tilde{\mathbf{f}}=\tilde{\mathbf{E}}_{k}\tilde{\mathbf{x}}$ for all $k\in\mathbf{\Omega}_{K}$. Therefore, any limit point of the sequence generated by Algorithm 1, $(\mathbf{F}^{(\infty)},\mathbf{X}^{(\infty)})$, is a stationary point of problem (8a).

### v3

@@ III-C3 Solving for Transmit Waveform
$\;\text{s.t. }\mathbf{X}\in\mathcal{S}_{\text{X}},$
whose closed-form solution is
@@ III-D Convergence Analysis
Any limit point of $\{(\mathbf{F}^{(\ell+1)},\mathbf{X}^{(\ell+1)})\}_{\ell\geq 0}$ generated by Algorithm 1 is a stationary point of problem (8).
@@ IV Numerical Results
In this section, we present simulation results to evaluate the performance of the proposed algorithms under various parameter configurations. The communication channel $\mathbf{H}$ is generated by following [32] with Extended Pedestrian A (EPA) fading profile [14]. The information symbol matrix $\mathbf{S}$ is modulated using a unit-power QPSK alphabet, with each entry in $\mathbf{S}$ having unit power. The desired radar receive spatial response, $\mathbf{R}_{\text{d}}$, is considered to be a $3$ dB beamwidth of $20^{\circ}$ focusing at $0^{\circ}$, generated by the algorithm in [40]. It is worth noting that the algorithm proposed in this paper can be directly extended to the Extremely large-scale MIMO (XL-MIMO) system [7] for generating unimodular low sidelobe level ISAC sequences.
@@
$K$
Maximum index of range bin of interest
6
@@
$P_{\text{F}}$
Power of the receive filter
64
@@ IV-A Convergence Performance
(b) The sum rate under SNR = 66 dB and beampattern MSE w.r.t $K$.
@@ IV-B Communication and Radar Performance under Different System Configurations
In this subsection, we evaluate the communication and radar performance of the proposed algorithm under different system configurations. Specifically, we examine the impact of the range sidelobe suppression area size determined by $K$ and the maximum sidelobe level constraints $\xi^{\prime}$ (dB).
Fig. 5 illustrates the impact of the sidelobe suppression area sizes determined by $K$ on the optimized results. As shown in Fig. 5LABEL:sub@fig:SL_K, even with the increment of $K$, the optimized sidelobe levels can still approach the predetermined threshold. In Fig. 5LABEL:sub@fig:rate_MSE_K, the achievable sum rate decreases, the beampattern MSE increases, and the mainlobe level decreases slightly (shown at the bottom-left of Fig. 5LABEL:sub@fig:SL_K) as $K$ grows, since a larger suppression area imposes more constraints, reducing the feasible region and degrading the performance of optimized results.
@@
(b) The sum rate under SNR = 66 dB and beampattern MSE w.r.t. PSLR constraints.
@@
Fig. 6 shows the impact of maximum sidelobe level constraints on the optimized results. As seen in Fig. 6LABEL:sub@fig:SL_PSLR, the sidelobe constraints can always be met. Lower sidelobe constraints require $\mathbf{X}$ and $\mathbf{F}$ to be less correlated, making the problem more complex and resulting in worse performance in both S&\&C. This can be observed in Fig. 6LABEL:sub@fig:rate_MSE_PSLR, in which a lower maximum sidelobe level constraint leads to a worse achievable sum rate and a larger beampattern MSE.
@@ IV-C Comparison with Existing SOTA Approaches
Figure 8: The MIMO radar image under the results of different algorithms when SNR = $15$ dB: (a) Original image of “T”. (b) The image formed by the proposed ALM algorithm. (c) The image formed by the Exact Design scheme. (d) The image formed by the fixed penalty ALM algorithm. (e) The image formed by the ADMM algorithm.
@@
In this subsection, we compare the radar and communication performance of the proposed approach with the modified work [29] and the ALM algorithm with a fixed penalty parameter. In [29], the joint design of the radar receive filter and transmit waveform focuses on PSL control for pure radar sensing. We adapt the algorithm in [29] by first solving $\min_{\mathbf{X}\in\mathcal{S}_{\text{X}}}\|\mathbf{Hx}-\mathbf{s}\|_{2}^{2}$ using the BSUM method and then following the approach in [29] to obtain the radar receive filter $\mathbf{F}$. We refer to the modified algorithm of [29] as the “Exact Design” scheme. Furthermore, we also implemented the ADMM algorithm and BSUM-based ALM algorithm with a fixed penalty parameter to show the superiority of the proposed adaptive penalty parameter scheme. We refer to them as “ADMM” and “Fixed Penalty” schemes, respectively. Without loss of generality, the penalty parameters for both the two algorithms are fixed to be $1$, and the total iterations of the “ADMM” algorithm are set the same as the proposed ALM algorithm. In the legends of the result figures, “Direc” refers to the receive beampattern focused at $30^{\circ}$, $0^{\circ}$, and $30^{\circ}$ with a $3$ dB beamwidth of $10^{\circ}$, generated by the algorithm in [40], and “Omni” denotes the omnidirectional desired beampattern, i.e., $\mathbf{R}_{\text{d}}=(P_{\text{X}}/N_{T})\mathbf{I}_{N_{T}}$.
@@
Fig. 7 presents the optimized results of different algorithms. In Fig. 7LABEL:sub@fig:rate_algs, the proposed ALM algorithm achieves the highest sum rate under the directional beampattern case. For an omnidirectional beam, we need to set a larger $\alpha$ to make the final beampattern closer to omnidirectional because an omnidirectional beam requires that $\mathbf{X}$ and $\mathbf{F}$ be orthogonal in each row and column, which significantly shrinks the feasible region. However, this degrades the minimization on $P_{\text{MUI}}$ and leads to a lower achievable rate. In the Exact Design scheme, although the transmit sequence is optimized solely to minimize the MUI at the receiver, it still has a lower spectral efficiency compared to the result of the proposed ALM algorithm. This suggests that directly minimizing $\|\mathbf{H}\mathbf{x}-\mathbf{s}\|_{2}^{2}$ is prone to suboptimal local minima. In contrast, our proposed algorithm demonstrates a strong ability to escape such local minima and attain better solutions, even under more stringent constraints. Other methods, due to the inappropriate penalty parameter, provide lower achievable rates. Fig. 7LABEL:sub@fig:bpMSE_algs shows that the Exact Design scheme nearly perfectly matches the desired beampatterns in both the directional and omnidirectional cases while the proposed inexact ALM algorithms have slight differences from the omnidirectional beampattern. The received beampatterns under the Fixed Penalty scheme and the ADMM scheme were far from the desired one because the inappropriate penalty parameter makes the algorithm focus too much on reducing the violation. Fig. 7LABEL:sub@fig:SL_algs shows the proposed ALM algorithm has the best sidelobe control performance.
Fig. 8 shows the MIMO radar images formed by the results of different algorithms [^3] We attempted to identify quantitative metrics for evaluating radar image quality, but to the best of our knowledge, no formal or widely accepted metric has been established. As in prior classical works [29, 41], visual inspection remains the standard evaluation approach, which we also adopt in this study.. The proposed inexact ALM algorithm achieved the best imaging performance. Although the Exact Design scheme perfectly meets the desired beampattern, it has a high sidelobe level, which results in severe interference caused by the clutter at the range bin of interest and, thus, a poorer MIMO radar image. Due to the inappropriate penalty parameter, the results under both the Fixed Penalty scheme and the ADMM scheme failed to produce radar images.
@@ Appendix D Proof of Theorem 2
The convergence proof consists of two main steps, where the first step shows the feasibility of the limit point of the sequence generated by Algorithm 1 and the second step shows that any limit point is a feasible stationary point of problem (8). We present these two steps in two subsections separately.
@@ D-A Feasibility of the Limit Point
D-B The Limit Point is a Stationary Feasible Solution of Problem (8)
Based on Definition 1 and Theorem 1, we know that as $\varepsilon^{(\ell)}\to 0$ and $\ell\to\infty$, any limit point generated by Algorithm 1 is a stationary point. In the previous subsection, we have shown that every such limit point is also a feasible solution to problem (LABEL:prob:FX_with_C). Hence, the pair $(\mathbf{F}^{(\infty)},\mathbf{X}^{(\infty)})$ constitutes a feasible stationary point of problem (LABEL:prob:FX_with_C). Furthermore, according to [38, Proposition 1], any feasible stationary point of the augmented Lagrangian problem is also a feasible stationary point of the original problem. Therefore, we conclude that every limit point generated by Algorithm 1 is a feasible stationary solution of the original problem (8).